OpenStudy (anonymous):

Find the slope-intercept form of the line whose slope is 2 and that passes through the point (-2,5)

OpenStudy (anonymous):

Use point-slope form. \[y - y_{1} = m(x - x _{1})\] \[y - 5 = 2(x + 2)\]

OpenStudy (anonymous):

How do you use this point slope form?

OpenStudy (anonymous):

x and y are the variables you use in slope intercept form, m is the slope, x1 and y1 are x and y coordinates of points on the line.

OpenStudy (anonymous):

How do I solve this problem? I am still lost right now.

OpenStudy (anonymous):

It was almost done at \[y - 5 = 2(x + 2)\] just add 5 to both sides and distribute the 2 over the (x + 2) and it will be in slope intercept form.

OpenStudy (anonymous):

I still do no understand how you get the answer. can you put it in a way I can understand it more easier?